PURE o2 diffuses through an aperture in 224 seconds,whereas mixture of o2 and anther gas containing 80% o2 diffuses from the same in 234 sec .what is the molecular weight of gas?

Grade:11

1 Answers

Arun
5 years ago

For gaseous mixture 80% O2, 20% of other gas m   

Average molecular weight Mm = 32 (for O2) × 80 + 20 × m / 100   

Now, for diffusion of gaseous mixture and pure O2   

rO2 / rm = √Mm / MO2 or VO2 / EO2 × Em / Vm = √Mm / MO2   

∴ 1 / 224 × 234 / 1 = √Mm / 32 

∴ Mm= 34.92  

∴ 32 × 80 + 20 × m / 100 = 34.92   

∴ m = 46.6

 

Regards

Arun